OpenStudy (anonymous):

What is the exact value of the expression √486−√24+√6? Simplify if possible.

OpenStudy (anonymous):

Do you remember how to simplify questions? I can bet that they will end up having like square roots once simplified...

OpenStudy (anonymous):

I mean simplify square roots

OpenStudy (anonymous):

well would √24 be 2√3?

OpenStudy (anonymous):

\[\sqrt{486} - \sqrt{24} + \sqrt{6}\implies \sqrt{81 \times 6} - \sqrt{4 \times 6} + \sqrt{6} \implies 9\sqrt{6} - 2\sqrt{6} + \sqrt{6}\]\[\implies \space ?\]

OpenStudy (anonymous):

I've honestly forgot. I've been working on this same test for 4&1/2 hours... my mind is wandering.

OpenStudy (anonymous):

8√6

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

thanks means a lot once again. 5 more questions to go...

OpenStudy (anonymous):

lol...don't worry...